OpenStudy (anonymous):

Two regular octagons have areas of 48 mm² and x² + 8x mm². What values of x make the octagons congruent?

OpenStudy (stephstar101):

x^2+8x=48 <- set the two areas equal to each other x^2+8x-48=0 <- subtract 48 from both sides (x-4)(x+12)=0 <-factor them out The last step is up to you! It'll give you the two final numbers, just solve for x!

OpenStudy (ganpat):

but the length cannot be negative.. So answer would be x= 4...

OpenStudy (anonymous):

my answers say 4 and 12 or 4 and -12

OpenStudy (stephstar101):

It's not asking for the length, just the values of x that make them congruent :D So any number can be x

OpenStudy (ganpat):

@stephstar101 : hmmm, yes any value can make it congruent.. @mr.perkins: so its 4 & -12..

OpenStudy (ganpat):

sorry, @mrs.perkins

OpenStudy (anonymous):

ok thank you